OpenStudy (anonymous):

Will Medal! Magnetic force decreases as distance from the magnet _____. increases decreases

OpenStudy (z4k4r1y4):

magnetic force is directly proportional to the magnetic field i.e. \[F_{B} \alpha B\]

OpenStudy (z4k4r1y4):

and magnetic field is inversely proportional to distance i.e. \[B \alpha \frac{ 1 }{ distance }\]

OpenStudy (z4k4r1y4):

so increasing the distance will reduce the magnetic field,B , and if B decreases the magnetic force,F , decreases.